Zachary Corporation paid one of its sales representatives $6,500 during the month of March. The rep is paid a base salary plu5
Answer #1

Total variable cost for March=(14*160)=$2240

Hence total fixed cost for March=6500-2240

=$4260

April May June July
Number of units 200 110 210 120
Total variable cost (14*200)=2800 (14*110)=1540 (14*210)=2940 (14*120)=1680
Total fixed cost 4260 4260 4260 4260
Total salary cost(Total fixed cost+Total variable cost) 7060 5800 7200 5940

NOTE:Total fixed costs and variable cost per unit do not change with change in units